# Enhancement of Magnon Transport by Superconductor Meissner Screening

### v1

@@ Abstract
Introduction.—Spin waves or their quanta, magnons, are collective excitations of ordered magnetic moments, which can strongly hybridize with other degrees of freedom for classical information transduction, quantum information processing, and long-range pure-spin transport [1, 32, 4, 5, 6, 7, 8, 3, 9]. In various magnonic devices, the intrinsic damping of spin waves is detrimental to their transport since it limits their possible propagation length, usually much smaller than photons and phonons. Ferrimagnetic insulator yttrium iron garnet (YIG) is currently the best material for magnon transport due to its record low damping [10, 11, 12, 13], long characteristic decay length of tens of micrometers [11, 13, 14], and narrowest known linewidth of ferromagnetic resonance, resulting in a lifetime of a few hundred nanoseconds [15, 16]. In spite of these merits, their dissipation parameters nowadays severely limit the appeal of employing magnons in the quantum information [8, 9, 17, 18, 19, 20, 21, 22] and quantum transport and appear to be challenging to be further improved via optimizing the crystal quality.
Superconducting devices are instruments implemented to avoid Ohmic dissipation and thereby employed on purpose in superconducting spintronics [25, 23, 24, 26, 27], quantum information [17, 18, 28, 19, 20, 29, 21, 30], and topological computation [31]. Useful ferromagnet and superconductor heterostructure are pursued to transport spin information without dissipation and realize new functionalities [25, 23, 24, 26, 27, 32]. Superconductors with thicknesses comparable to London’s penetration depth allow the penetration of stray field emitted by spin waves, providing the electromotive force that drives the supercurrent and in turn, causes the Ostered magnetic field that shifts the spin-wave frequency by many gigahertz [33, 34, 36, 37, 38, 39, 35, 40, 46, 41, 43, 42, 44, 45]. However, the role of dissipationless supercurrent on magnon transport via this “gating” effect has not yet been explored in previous theories [33, 34, 36, 37, 38, 39, 35, 40, 46, 41, 43, 42, 44, 45], raising the question that whether the supercurrent can “shortcut” the magnon energy flow as in electronics.
@@
Figure 1: Inductive coupling between magnetization dynamics and Cooper-pair supercurrent in FI-SC heterostructure. An in-plane external magnetic field ${\bf H}_{0}$ biases ${\bf M}_{0}$ to the $\hat{\bf z}$-direction. Spin waves propagating normally to the saturated magnetization ${\bf M}_{0}$ radiate the electric fields along ${\bf M}_{0}$. An inserted thin insulator spacer is assumed to suppress the exchange proximity effect.
We find that the electromotive force and supercurrent induced by the stray magnetic fields emitted by the magnetization dynamics can reflect, focus, and enhance the stray field inside the FI, such that the group velocity of spin waves is significantly enhanced, while the net damping is not suppressed. We predict that by changing the temperature across the superconducting transition temperature (e.g., 5-10 K for NbN [47, 48, 49]), the magnon transport can be unidirectionally enhanced by up to $450\%$ for YIG film capped by superconducting NbN film, rendering a decay length exceeding millimeters, well beyond the present records [10, 11, 12, 13, 14]. Since quantum transport and quantum information are performed at cryogenic temperatures, our findings may augment the performance of magnons in these fields.
Reflection of magnetization radiation by proximity superconductors.—Here we analytically calculate the electric and magnetic fields radiated by the magnetization fluctuation in the FI-SC heterostructure by solving Maxwell’s equations and matching the boundary conditions at interfaces between the FI, SC, and vacuum. As shown in Fig. 1, a static in-plane magnetic field ${\bf H}_{0}$ orients the saturated magnetization ${\bf M}_{0}$ to the $\hat{\bf z}$-direction and the thickness of FI and SC are $2d_{F}$ and $d_{S}$, respectively. For the transverse magnetization fluctuation $M_{x,y}({\bf r},t)={\cal M}_{x,y}(x)e^{i{\bf k}\cdot{\boldsymbol{\rho}}}e^{-i\omega t}$ of frequency $\omega$ with amplitudes ${\cal M}_{x,y}$ and wave vector $\mathbf{k}=(0,k_{y},k_{z})$ propagating in the film plane ${\boldsymbol{\rho}}=y\hat{\bf y}+z\hat{\bf z}$, ${\cal M}_{x,y}$ is nearly uniform across the film normal direction (i.e., the $\hat{\bf x}$-direction) when the film is sufficiently thin and the wavelength of spin waves ($\gg 2d_{F}$) is long [35, 10, 11, 12, 13, 14]. Such magnetic fluctuation radiates the electric field ${\bf E}({\bf r},t)$ via the magnetization current ${\bf J}_{M}({\bf r},t)=\nabla\times{\bf M}({\bf r},t)$. According to Maxwell’s equations [50], the radiation obeys $\nabla^{2}\mathbf{E}({\bf r},t)+k^{2}_{0}\mathbf{E}({\bf r},t)=-i\omega\mu_{0}{\bf J}_{M}({\bf r},t)$, where $k_{0}=\omega\sqrt{\mu_{0}\varepsilon_{0}}$ with $\mu_{0}$ and $\varepsilon_{0}$ being the vacuum permeability and permittivity, respectively.
@@
where ${\cal A}_{k}=\sqrt{k_{0}^{2}-k^{2}}\rightarrow ik$, $\alpha_{k}(x)=e^{i{\cal A}_{k}\left(x+d_{F}\right)}-e^{i{\cal A}_{k}\left(x-d_{F}\right)}$, $\delta_{\pm}(k,x)=e^{-i{\cal A}_{k}\left(x-d_{F}\right)}\pm e^{i{\cal A}_{k}\left(x+d_{F}\right)}$, and $\beta_{\bf k}=-ia_{\bf k}k_{y}/{\cal A}_{k}\rightarrow-a_{\bf k}k_{y}/k$. When the spin waves propagate normally to ${\bf M}_{0}$, the electric field is linearly polarized along the saturated magnetization $\hat{\bf z}$-direction since ${\cal E}^{(0)}_{x/y}$ vanishes when $k_{z}=0$. Particularly, for circularly polarized spin waves with $a_{\bf k}=1$, $\beta_{\bf k}\sim-a_{\bf k}k_{y}/k=-\text{sgn}{(k_{y})}$ depends on the propagation direction: above (below) the ferromagnetic film, the electric field exists only when $k_{y}>0$ ($k_{y}<0$), governed by the attenuation length $|k_{y}|^{-1}$, but vanishes when $k_{y}<0$ ($k_{y}>0$).
We then demonstrate that the radiated electric (and magnetic) field by a single FI is strongly reflected, focused, and enhanced by the adjacent SCs. Inside the SCs, the electric field is screened via $\nabla^{2}\mathbf{E}({\bf r},t)+k_{s}^{2}\mathbf{E}({\bf r},t)=0,$ where $k_{s}=\sqrt{\omega^{2}\mu_{0}\varepsilon_{0}-{1}/{\lambda_{L}^{2}}}$ is the electromagnetic wave number in the SC, which is approximately purely imaginary and governed by the inverse London’s penetration depth $\lambda_{L}$ [50]. Therefore, the electric field is screened by the supercurrent and penetrates only within $\lambda_{L}$. The temperature $T$ affects the penetration according to $\lambda_{L}({T})=\lambda_{L}(T=0~{}{\rm K})/\sqrt{1-(T/T_{c})^{4}}$ [51], which is large when $T$ approaches the superconducting transition temperature $T_{c}$.
@@
The boundary condition governs the reflection and transmission of electric fields at interfaces. The in-plane ${\bf E}_{\parallel}({\bf r},t)$ is continuous at interfaces [50]. For the normal component $E_{x}$, there is no charge accumulation inside SCs at microwave frequency, such that $\nabla\cdot{\bf E}({\bf r},t)=0$ in it. On the other hand, $\partial_{\parallel}{\bf E}_{\parallel}({\bf r},t)$ is continuous, and thereby, combining with $\nabla\cdot{\bf E}({\bf r},t)=0$, $\partial_{x}E_{x}({\bf r},t)$ is continuous at interfaces. Finally, the current $J_{x}({\bf r},t)+\varepsilon_{0}\partial_{t}E_{x}({\bf r},t)$ is continuous at interfaces due to the charge conservation. Carefully matching the boundary conditions refine the previous theoretical treatments [35, 33], by which we obtain the electric field for all wave vectors $\bf k$, referring to the Supplemental Material for details [52].
@@
Here, we focus on the Damon-Eshbach (DE) configuration with $k_{z}=0$ since these spin waves are modulated most strongly by SCs and were recently measured [35]. In this case, as that radiated by a single FI (1) only the electric field $E_{z}\hat{\bf z}\parallel{\bf M}_{0}$ exists, as addressed in Fig. 1. The electric field emitted by a single FI (1) is reflected by the SC according to the wave-number dependent reflection coefficient
${\cal R}_{k}=({\cal A}_{k}-{\cal B}_{k})/({\cal A}_{k}+{\cal B}_{k}),$ (2)
where ${\cal B}_{k}=\sqrt{k_{s}^{2}-k^{2}}$ and $k=|k_{y}|$ in the DE configuration, such that inside the FI
@@
For a large $k\gg\lambda_{L}^{-1}$, the electric field rapidly decays within the decay length $k^{-1}\ll\lambda_{L}$, such that $|{\cal A}_{k}|\sim|{\cal B}_{k}|$, implying the vanished reflection (2) ${\cal R}_{k}\sim({\cal A}_{k}-{\cal A}_{k})/(2{\cal A}_{k})\rightarrow 0$, i.e., the SC hardly affects the spin waves. On the other hand, when $k\ll\lambda_{L}^{-1}$, $|{\cal A}_{k}|\ll|{\cal B}_{k}|$ and ${\cal R}_{k}\sim-{\cal B}_{k}/{\cal B}_{k}\rightarrow-1$, such that the electric field is totally reflected by the SC with a $\pi$ phase shift, resulting in no net supercurrents driven inside the SC, i.e., the magnon and Cooper-pair supercurrent decouples. Thereby, only the spin waves with $k\sim\lambda_{L}^{-1}$ are most strongly modulated by the SCs.
The electric field (3) induces the magnetic field ${\bf H}_{d}({\bf r},t)$ according to Faraday’s Law $i\omega\mu_{0}[{\bf H}_{d}({\bf r},t)+{\bf M}({\bf r},t)]=\nabla\times{\bf E}({\bf r},t)$ [50]. As addressed before, the magnetic fluctuation is nearly uniform across the thin FI, and we thereby take the spatial average for the stray field along the $\hat{\bf x}$-direction:
@@
where the form factor $P_{k}=\sinh(kd_{F})/(kd_{F})$. As in the electric field (3), the dipolar field is also expressed as the one radiated by a single FI superposed by that reflected by the SCs. When $k_{y}\rightarrow 0$ or ${\cal R}_{k}\rightarrow 0$, the influence of the SC vanishes, and the dipolar field recovers to that of a single FI, which, however, is enhanced when ${\cal R}_{k}\rightarrow-1$ assuming $a_{k_{y}}\sim 1$. The dimensionless factors $\kappa_{y}(k_{y})<0$ and $\kappa_{x}(k_{y})<0$ are generally enhanced in magnitude by the SC. The chirality of the dipolar field is now clear: for circularly polarized spin waves with $a_{k_{y}}\rightarrow 1$, $[1+a_{k_{y}}\text{sgn}(k_{y})]/2\rightarrow 0$ and $[1+a^{-1}_{k_{y}}\text{sgn}(k_{y})]/2\rightarrow 0$ when $k_{y}<0$, but becomes unity when $k_{y}>0$.
We now illustrate the effect of SCs on the stray field by comparing it with that of an isolated FI in Fig. 2(a) and (b). We take $d_{F}=50$ nm, $d_{S}=300$ nm, and London’s penetration depth $\lambda_{L}=80$ nm at $T=1.2~{}\text{K}\sim 0.1T_{c}$ for NbN [47, 48, 49]. For the spin wave with $k_{y}=4$ $\mu{\rm m}^{-1}$, without gating its stray field mainly exists above the ferromagnetic film $x>0$ [Fig. 2(a)]. When in proximity to the SC, the magnetic field at $x>0$ is reflected and decays faster inside the SC, which is then enhanced inside the ferromagnetic film. In the present case, the enhancement is up to 30% compared to that with a single FI. By contrast, when $k_{y}=-4$ $\mu{\rm m}^{-1}$, as shown in Fig. 2(b), the emitted magnetic field without gating mainly exists below the ferromagnetic film $x<0$, thereby is hardly affected by the SC fabricated at $x>0$.
@@
where $H_{k_{y}}=H_{0}+\alpha_{\rm ex}k_{y}^{2}M_{0}$. Equation (5c) provides the self-consistent equation for solving the ellipticity, with which we numerically find the dispersion (5a) and inverse lifetime (5b). When $k_{y}>0$, $\kappa_{x}(k_{y})<0$ and $\kappa_{y}(k_{y})<0$ are both enhanced in magnitude by the adjacent SC, such that both the frequency (5a) and the inverse lifetime (5b) are enhanced. It is then not clear whether the transport can be definitely enhanced or not.
@@
where $\tilde{k}_{y}=k_{\Omega}+i\tilde{k}_{\Omega}$ is the positive root of $\tilde{\omega}(\tilde{k}_{y})=\Omega$ and $v_{\bf k}=\partial\omega({\bf k})/\partial{\bf k}$ is the group velocity of spin waves. $\lambda=1/\tilde{k}_{\Omega}$ governs the decay length of the excited magnetization.
We find $\tilde{k}_{\Omega}$ as follows. Since $\tilde{\omega}({k_{y}})$ is solved from the linearized LLG equation, $\tilde{\omega}({k_{y}})=\Omega$ obeys
$\Omega^{2}=$ $\left[\mu_{0}\gamma(H_{k_{y}}-M_{0}\kappa_{y}({k_{y}}))-i\Omega\alpha_{G}\right]$
$\times$ $\left[(\mu_{0}\gamma(H_{k_{y}}-M_{0}\kappa_{x}({k_{y}}))-i\Omega\alpha_{G}\right]$
$=$ $\omega^{2}(k_{y})-2i\Omega\Gamma(k_{y}).$ (9)
The solution $\tilde{k}_{y}=k_{\Omega}+i\tilde{k}_{\Omega}$ is a complex number containing a small imaginary component $\tilde{k}_{\Omega}\ll k_{\Omega}$, which allows us to expand $\omega(\tilde{k}_{y})$ to the leading order as
@@
Figure 3: Enhancement of magnon transport in FI by an adjacent SC in the DE configuration ${\bf k}\perp{\bf M}_{0}$. (a), (b), and (c) plot the frequency shift, enhanced group velocity, and lifetime of magnons when the temperature is decreased from $T=T_{c}$ to $0.1T_{c}$. These modulations by the adjacent SC lead to a significant enhancement of magnon transport as demonstrated in the dependencies of the characteristic decay length on the space [(d)], excited frequency $\Omega$ [(e)], and temperatures [(f)].

### v2

# Giant enhancement of magnon transport by superconductor Meissner screening

@@ Abstract
Introduction.—Spin waves or their quanta, magnons, are collective excitations of ordered magnetic moments, which can strongly hybridize with other degrees of freedom for classical information transduction, quantum information processing, and long-range pure-spin transport [1, 2, 4, 5, 6, 7, 8, 3, 9]. In various magnonic devices, the intrinsic damping of spin waves is detrimental to their transport since it limits their possible propagation length, usually much smaller than photons and phonons. Ferrimagnetic insulator yttrium iron garnet (YIG) is currently the best material for magnon transport due to its record low damping [10, 11, 12, 13], long characteristic decay length of tens of micrometers [11, 13, 14], and narrowest known linewidth of ferromagnetic resonance, resulting in a lifetime of a few hundred nanoseconds [16, 15]. The coherence time must be pushed to its limits to realize, implement, and assess the quantum states of magnons [8, 9, 17, 18, 20, 21, 22, 19]. Despite these merits, their dissipation parameters severely limit the appeal of employing magnons in quantum transport, which appears challenging to improve further via optimizing crystal quality.
Superconducting devices are instruments implemented to avoid Ohmic dissipation and employed on purpose in superconducting spintronics [25, 23, 24, 26, 27], quantum information [17, 18, 28, 20, 21, 29, 22, 30], and topological computation [31]. Useful ferromagnet and superconductor heterostructure are pursued to transport spin information without dissipation and realize new functionalities [25, 23, 24, 26, 27, 32, 33]. Superconductors with thicknesses comparable to London’s penetration depth allow penetration of stray field emitted by spin waves, providing the electromotive force that drives the supercurrent and, in turn, causes the Ostered field that shifts the spin-wave frequency by many gigahertz [42, 43, 34, 35, 37, 39, 45, 47, 46, 36, 40, 38, 41, 44]. However, the role of dissipationless supercurrent on magnon transport via this “gating” effect has not yet been explored in previous theories [42, 43, 34, 35, 37, 39, 45, 47, 46, 36, 40, 38, 41, 44], raising the question that whether the supercurrent can “shortcut” the magnon energy flow as in electronics.
@@
Figure 1: Inductive coupling between magnetization dynamics and Cooper-pair supercurrent in FI-SC heterostructure. An in-plane external magnetic field ${\bf H}_{0}$ biases ${\bf M}_{0}$ to the $\hat{\bf z}$-direction. Spin waves propagating normally to the saturated magnetization ${\bf M}_{0}$ radiate the electric fields along ${\bf M}_{0}$.
We find that the electromotive force and supercurrent induced by the stray magnetic fields emitted by the magnetization dynamics can reflect, focus, and enhance the stray field inside the FI, such that the group velocity of spin waves is significantly enhanced while the net damping is not suppressed. We predict that by changing the temperature across the superconducting transition temperature (e.g., 5-10 K for NbN [48, 49, 50]), the magnon transport can be unidirectionally enhanced by up to $450\%$ for YIG film capped by superconducting NbN film, rendering a decay length exceeding millimeters, well beyond the present records [10, 11, 12, 13, 14]. Since quantum transport and quantum information are performed at cryogenic temperatures, our findings may augment the performance of magnons in these fields.
Reflection of magnetization radiation by proximity superconductors.—Here we analytically calculate the electric and magnetic fields radiated by the magnetization fluctuation in the FI-SC heterostructure by solving Maxwell’s equations and matching the boundary conditions at interfaces between the FI, SC, and vacuum. The exchange proximity effect between SCs and FIs can be intentionally suppressed by inserting an insulator spacer such as Al2O3 [51], leaving the long-range dipolar interaction unaffected. As shown in Fig. 1, a static in-plane magnetic field ${\bf H}_{0}$ orients the saturated magnetization ${\bf M}_{0}$ to the $\hat{\bf z}$-direction and the thickness of FI and SC are $2d_{F}$ and $d_{S}$, respectively. For the transverse magnetization fluctuation $M_{x,y}({\bf r},t)={\cal M}_{x,y}(x)e^{i{\bf k}\cdot{\boldsymbol{\rho}}}e^{-i\omega t}$ of frequency $\omega$ with amplitudes ${\cal M}_{x,y}$ and wave vector $\mathbf{k}=(0,k_{y},k_{z})$ propagating in the film plane ${\boldsymbol{\rho}}=y\hat{\bf y}+z\hat{\bf z}$, ${\cal M}_{x,y}$ is nearly uniform across the film normal direction (i.e., the $\hat{\bf x}$-direction) when the film is sufficiently thin and the wavelength of spin waves ($\gg 2d_{F}$) is long [45, 10, 11, 12, 13, 14]. Such magnetic fluctuation radiates the electric field ${\bf E}({\bf r},t)$ via the magnetization current ${\bf J}_{M}({\bf r},t)=\nabla\times{\bf M}({\bf r},t)$. According to Maxwell’s equations [52], the radiation obeys $\partial_{x}^{2}\mathbf{E}({\bf r},t)+(k^{2}_{0}-k^{2})\mathbf{E}({\bf r},t)=-i\omega\mu_{0}{\bf J}_{M}({\bf r},t)$, where $k_{0}=\omega\sqrt{\mu_{0}\varepsilon_{0}\varepsilon_{r}}$ with $\mu_{0}$, $\varepsilon_{0}$, and $\varepsilon_{r}$ being the vacuum permeability, permittivity, and relative dielectric constant, respectively. $k_{0}\sim 0.01~{}\mu{\rm m}^{-1}$ is tiny even for large $\varepsilon_{r}\sim 10$ for YIG and $\omega\sim 100$ GHz.
@@
where ${\cal A}_{k}=\sqrt{k_{0}^{2}-k^{2}}\rightarrow ik$, $\alpha_{k}(x)=e^{i{\cal A}_{k}\left(x+d_{F}\right)}-e^{i{\cal A}_{k}\left(x-d_{F}\right)}$, $\delta_{\pm}(k,x)=e^{-i{\cal A}_{k}\left(x-d_{F}\right)}\pm e^{i{\cal A}_{k}\left(x+d_{F}\right)}$, and $\beta_{\bf k}=-ia_{\bf k}k_{y}/{\cal A}_{k}\rightarrow-a_{\bf k}k_{y}/k$. When spin waves propagate normally to ${\bf M}_{0}$, the electric field is linearly polarized along the saturated magnetization $\hat{\bf z}$-direction since ${\cal E}^{(0)}_{x/y}$ vanishes when $k_{z}=0$. Particularly, for circularly polarized spin waves with $a_{\bf k}=1$, $\beta_{\bf k}\sim-a_{\bf k}k_{y}/k=-\text{sgn}{(k_{y})}$ depends on the propagation direction: above (below) the ferromagnetic film, the electric field exists only when $k_{y}>0$ ($k_{y}<0$), governed by the attenuation length $|k_{y}|^{-1}$, but vanishes when $k_{y}<0$ ($k_{y}>0$).
This chiral electric field arises from the interference of surface and bulk contributions in the magnetization current ${\bf J}_{M}(x)=\nabla\times{\bf M}=[\delta(x+d_{F})-\delta(x-d_{F})-k_{y}]M_{y}\hat{\bf z}$. The volume current $-k_{y}M_{y}\hat{\bf z}$ is proportional to the propagation direction $k_{y}$, which radiates the electric field $E_{z}$ decaying away from the surface. The direction $k_{y}$ determines the direction of $E_{z}$, as illustrated in Fig. 2(a) and (b) for $k_{y}>0$ and $k_{y}<0$. On the other hand, the surface current $\delta(x+d_{F})M_{y}-\delta(x-d_{F})M_{y}$ is located at the upper and lower surfaces flowing in opposite directions, which radiates the electric field polarized in opposite directions above and below the magnetic film, as shown in Fig. 2(c). Thereby, the net electric field, contributed by both the volume and surface magnetization currents, is suppressed (enhanced) below the film but is enhanced (suppressed) above the film when $k_{y}>0$ ($k_{y}<0$), exhibiting the unidirectionality.
Figure 2: Snapshot of electric-field distribution radiated by the volume magnetization current when $k_{y}>0$ [(a)], when $k_{y}<0$ [(b)], and that by the surface magnetization current [(c)].
We then demonstrate that the radiated electric (and magnetic) field by a single FI is strongly reflected, focused, and enhanced by the adjacent SCs. Inside the SCs, the electric field generates both normal current ${\bf J}_{n}$ and supercurrent ${\bf J}_{s}$ [53, 54]. The temperature $T$ affects them via the superfluid density $\rho_{s}(T)=\rho_{e}(1-(T/T_{c})^{4})$ and normal fluid density $\rho_{n}(T)=\rho_{e}(T/T_{c})^{4}$ when $T<T_{c}$. $\rho_{s}+\rho_{n}$ equals the electron density $\rho_{e}$. Then the conductivity at low frequencies
$\tilde{\sigma}(\omega)\approx\dfrac{\rho_{n}(T)e^{2}\tilde{\tau}}{m_{e}}+i\dfrac{\rho_{s}(T)e^{2}}{m_{e}\omega}=\sigma_{n}(T)+i\frac{1}{\omega\mu_{0}\lambda^{2}_{L}(T)},$
where $\tilde{\tau}$ is the relaxation time of electrons and $m_{e}$ is the electron mass, $\sigma_{n}(T)=\rho_{n}(T)e^{2}\tilde{\tau}/m_{e}$ is the conductivity of normal fluid and $\lambda_{L}(T)=\sqrt{m_{e}/(\mu_{0}\rho_{s}(T)e^{2})}$ [55] is London’s penetration depth. The total electric current ${\bf J}=\tilde{\sigma}{\bf E}={\bf J}_{n}+{\bf J}_{s}$ screens the electric field via $\nabla^{2}\mathbf{E}({\bf r},t)+k_{s}^{2}\mathbf{E}({\bf r},t)=0,$ where $k_{s}=\sqrt{\omega^{2}\mu_{0}\varepsilon_{0}-{1}/{\lambda_{L}(T)^{2}}+i\omega\mu_{0}\sigma_{n}(T)}$ is the electromagnetic wave number in the SC, which is complex governed by $\lambda_{L}(T)$ and $\sigma_{n}(T)$. Although the normal fluid is included in the formalism, it plays a minor role when $T<0.95T_{c}$ since $\sigma_{n}(T)\propto(T/T_{c})^{4}$ rapidly decays when decreasing the temperature, referring to the Supplemental Material (SM) for details [56].
@@
The boundary condition governs the reflection and transmission of electric fields at interfaces. The in-plane ${\bf E}_{\parallel}({\bf r},t)$ is continuous at interfaces [52]. For the normal component $E_{x}$, there is no charge accumulation inside SCs at microwave frequency, such that $\nabla\cdot{\bf E}({\bf r},t)=0$ in it. On the other hand, $\partial_{\parallel}{\bf E}_{\parallel}({\bf r},t)$ is continuous, and thereby, combining with $\nabla\cdot{\bf E}({\bf r},t)=0$, $\partial_{x}E_{x}({\bf r},t)$ is continuous at interfaces. Finally, the current $J_{x}({\bf r},t)+\varepsilon_{0}\partial_{t}E_{x}({\bf r},t)$ is continuous at interfaces due to the charge conservation. Carefully matching the boundary conditions go beyond the previous theoretical treatments [45, 42], by which we obtain the magnetization radiation for all wave vectors $\bf k$, referring to the SM for details [56].
@@
Here, we focus on the Damon-Eshbach (DE) configuration with $k_{z}=0$ since these spin waves are modulated most strongly by SCs and were recently measured [45]. In this case, as that radiated by a single FI (1) only the electric field $E_{z}\hat{\bf z}\parallel{\bf M}_{0}$ exists, as addressed in Fig. 1. The electric field emitted by a single FI (1) is reflected by the SC according to the wave-number dependent reflection coefficient ${\cal R}_{k}=({\cal A}_{k}-{\cal B}_{k})/({\cal A}_{k}+{\cal B}_{k})$, where ${\cal B}_{k}=\sqrt{k_{s}^{2}-k^{2}}$ and $k=|k_{y}|$ in the DE configuration, such that inside the FI
@@
For a large $k\gg\lambda_{L}^{-1}$, the electric field rapidly decays within the decay length $k^{-1}\ll\lambda_{L}$, such that $|{\cal A}_{k}|\sim|{\cal B}_{k}|$, implying vanished reflection ${\cal R}_{k}\sim({\cal A}_{k}-{\cal A}_{k})/(2{\cal A}_{k})\rightarrow 0$, i.e., the SC hardly affects the spin waves. On the other hand, when $k=0$, $|{\cal A}_{k}|\ll|{\cal B}_{k}|$ and ${\cal R}_{k}\sim-{\cal B}_{k}/{\cal B}_{k}=-1$, such that the electric field is totally reflected by the SC with a $\pi$ phase shift, resulting in no net supercurrents driven inside the SC, i.e., the magnon and Cooper-pair supercurrent decouples. However, the SC modulates the spin waves even for small $k$ since ${\cal R}_{k}$ rapidly departs from $-1$. The spin waves with $k\sim\lambda_{L}^{-1}$ are most strongly modulated by SCs.
The electric field (2) induces the magnetic field ${\bf H}_{d}({\bf r},t)$ according to Faraday’s Law $i\omega\mu_{0}[{\bf H}_{d}({\bf r},t)+{\bf M}({\bf r},t)]=\nabla\times{\bf E}({\bf r},t)$ [52]. Inside the FI, the magnetic stray field ${\bf H}_{d}({\bf r},t)$ decays with the decay length $\sim 1/|k|$, while the attenuation length $\sim(k^{2}+1/\lambda_{L}^{2})^{-1/2}$ in SCs is much shorter when $|k|<\lambda_{L}^{-1}$. Thereby, even for thin SCs of tens of nanometers, the decay is significant, while it is not equally important in the thin FI of similar thickness. We thereby take the spatial average for the stray field in the FI along the normal $\hat{\bf x}$-direction:
@@
where the form factor $P_{k}=\sinh(kd_{F})/(kd_{F})$. As in the electric field (2), the dipolar field is also expressed as the one radiated by a single FI superposed by that reflected by the SCs. When $k_{y}=0$ or ${\cal R}_{k}\rightarrow 0$, the influence of the SC vanishes, and the dipolar field recovers to that of a single FI, which, however, is enhanced when ${\cal R}_{k}\rightarrow-1$ assuming $a_{k_{y}}\sim 1$. The dimensionless factors $\kappa_{y}(k_{y})<0$ and $\kappa_{x}(k_{y})<0$ are generally enhanced in magnitude by the SC. The chirality of the dipolar field is now clear: for circularly polarized spin waves with $a_{k_{y}}\rightarrow 1$, $[1+a_{k_{y}}\text{sgn}(k_{y})]/2\rightarrow 0$ and $[1+a^{-1}_{k_{y}}\text{sgn}(k_{y})]/2\rightarrow 0$ when $k_{y}<0$, but becomes unity when $k_{y}>0$.
We now illustrate the effect of SCs on the stray field by comparing it with that of an isolated FI in Fig. 3(a) and (b). We take $d_{F}=50$ nm, $d_{S}=300$ nm, and London’s penetration depth $\lambda_{L}=80$ nm at $T=1.2~{}\text{K}\sim 0.1T_{c}$ for NbN [48, 49, 50]. For the spin wave with $k_{y}=4$ $\mu{\rm m}^{-1}$, without gating its stray field mainly exists above the ferromagnetic film $x>0$ [Fig. 3(a)]. When in proximity to the SC, the magnetic field at $x>0$ is reflected and decays faster inside the SC, which is then enhanced inside the ferromagnetic film. By contrast, when $k_{y}=-4$ $\mu{\rm m}^{-1}$, as shown in Fig. 3(b), the emitted magnetic field without gating mainly exists below the ferromagnetic film $x<0$, thereby is hardly affected by the SC fabricated at $x>0$.
@@
where $H_{k_{y}}=H_{0}+\alpha_{\rm ex}k_{y}^{2}M_{0}$. Equation (4c) provides the self-consistent equation for solving the ellipticity, with which we numerically find the dispersion (4a) and inverse lifetime (4b). When $k_{y}>0$, $\kappa_{x}(k_{y})<0$ and $\kappa_{y}(k_{y})<0$ are both enhanced in magnitude by the adjacent SC, such that both the frequency (4a) and the inverse lifetime (4b) are enhanced. It is then unclear whether the transport can be definitely enhanced.
@@
where $\tilde{k}_{y}=k_{\Omega}+i\tilde{k}_{\Omega}$ is the positive root of $\tilde{\omega}(\tilde{k}_{y})=\Omega$, and $v_{\bf k}=\partial\omega({\bf k})/\partial{\bf k}$ is the group velocity of spin waves. $\lambda=1/\tilde{k}_{\Omega}$ governs the decay length of excited magnetization.
We find $\tilde{k}_{\Omega}$ as follows. From the linearized LLG equation
$-i\omega M_{x}+\mu_{0}\gamma H_{k_{y}}M_{y}=\mu_{0}\gamma M_{0}\kappa_{y}(k_{y})M_{y}+i\alpha_{G}\omega M_{y},$
$i\omega M_{y}+\mu_{0}\gamma H_{k_{y}}M_{x}=\mu_{0}\gamma M_{0}\kappa_{x}(k_{y})M_{x}+i\alpha_{G}\omega M_{x},$
the eigenfrequency $\tilde{\omega}(k_{y})$ is governed by the characteristic equation
$\tilde{\omega}^{2}(k_{y})$ $=\left[\mu_{0}\gamma(H_{k_{y}}-M_{0}\kappa_{y}({k_{y}}))-i\tilde{\omega}(k_{y})\alpha_{G}\right]$
×$\times\left[\mu_{0}\gamma(H_{k_{y}}-M_{0}\kappa_{x}({k_{y}}))-i\tilde{\omega}(k_{y})\alpha_{G}\right].$ (8)
$\tilde{\omega}(\tilde{k}_{y})=\Omega$ leads to
$\Omega^{2}$ $=\left[\mu_{0}\gamma(H_{\tilde{k}_{y}}-M_{0}\kappa_{y}({\tilde{k}_{y}}))-i\Omega\alpha_{G}\right]$
×$\times\left[\mu_{0}\gamma(H_{\tilde{k}_{y}}-M_{0}\kappa_{x}({\tilde{k}_{y}}))-i\Omega\alpha_{G}\right]$
$=\omega^{2}(\tilde{k}_{y})-2i\Omega\Gamma(\tilde{k}_{y})-\Omega^{2}\alpha_{G}^{2}$
$\approx\omega^{2}(\tilde{k}_{y})-2i\Omega\Gamma(\tilde{k}_{y}),$ (9)
where the quadratic term $\Omega^{2}\alpha_{G}^{2}$ is disregarded since $\alpha_{G}\ll 1$. The solution $\tilde{k}_{y}=k_{\Omega}+i\tilde{k}_{\Omega}$ is a complex number containing a small imaginary component $\tilde{k}_{\Omega}\ll k_{\Omega}$, such that $\Omega\Gamma(\tilde{k}_{y})\sim\Omega\Gamma({k}_{\Omega})$, and we are allowed to expand $\omega(\tilde{k}_{y})$ to the leading order as
@@
Figure 4: Enhancement of magnon transport in FI by an adjacent SC in the DE configuration ${\bf k}\perp{\bf M}_{0}$. (a), (b), and (c) plot the frequency shift, enhanced group velocity, and lifetime of magnons when the temperature $T=0.1T_{c}$ compared with that without SC gating. These modulations by the adjacent SC lead to a significant enhancement of magnon transport as demonstrated in the dependencies of the excited frequency $\Omega$ [(d)], characteristic decay length in the space [(e)], and temperatures [(f)].